# Unlocking Doping Effects on Altermagnetism in MnTe

### v1

@@ IV Summary
For paired non-magnetic substitutions, the supercell produces 120 possible configurations, and based on their space group symmetries, they are segregated into five different families. We discovered that perfect altermagnetism persists in three families covering $\approx$ 46% of all possible configurations, while in the remaining two, the opposite-spin sublattices are not rotational-symmetry related and thus do not qualify as ideal altermagnets. However, their band structures retain the key altermagnetic features, including momentum-dependent spin splitting and nearly compensated net magnetization. For isovalent doping, deviations from perfect altermagnetism are minimal, whereas hole and electron doping produce more pronounced deviations, but still result in near-zero magnetization due to alternating spin dominance across the Brillouin zone. We have termed this behavior as quasi-altermagnetism, characterized by significant momentum-dependent spin splitting with a small residual moment. Moreover, through model Hamiltonians, we investigated the microscopic origin of the quasi-altermagnetism and identify the perturbations that lead to the transition from ideal altermagnetism to quasi-altermagnetism.
@@ Acknowledgments
N.D. would like to acknowledge the IoE-IISc postdoctoral fellowship. A.M. thanks MoE India for the PMRF fellowship. M.A.R. is supported by a graduate fellowship of the Indian Institute of Science. A.D. is thankful for the graduate fellowship of the IIT Madras. A.N. acknowledges support from Anusandhan National Research Foundation (ANRF) through grant number CRG/2023/000114.
@@ Appendix B Symmetry operations connecting opposite spin sublattices in doped MnTe structures
For the single dopant scenario, the system consistently exhibits altermagnetic ordering. In contrast, the paired dopant scenario leads to several distinct configurations [see Fig. 3(e)], with 46.67% maintaining altermagnetism, while the rest do not. Table 1 summarizes the symmetry operations connecting the opposite spin sublattices and corresponding magnetic classes for all these doped systems. For MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211), $Fm^{\prime}m^{\prime}2$ (No.42.222), and $Pm^{\prime}m^{\prime}2$ (No.25.60) the opposite spin sub-lattices are connected by symmetries other than translation or inversion, therefore being classified as altermagnets. On the other hand, for $P\bar{3}m^{\prime}1$ (No.164.89) and $C2^{\prime}/m^{\prime}2$ (No.12.62), which we refer to as quasi-altermagnets, there exist no symmetry operations that connect the up and down spin sublattices. This absence results in a lack of compensation between the spin channels, thereby lifting the degeneracy between the up and down spin bands throughout the Brillouin zone.

### v2

@@ IV Summary
For paired non-magnetic substitutions, the supercell produces 120 possible configurations, and based on their space group symmetries, they are segregated into five different families. We discovered that perfect altermagnetism persists in three families covering $\approx$ 46% of all possible configurations, while in the remaining two, the opposite-spin sublattices are not rotational-symmetry related and thus do not qualify as ideal altermagnets. However, their band structures retain the key altermagnetic features, including momentum-dependent spin splitting and nearly compensated net magnetization. For isovalent doping, deviations from perfect altermagnetism are minimal, whereas hole and electron doping produce more pronounced deviations, but still result in near-zero magnetization due to alternating spin dominance across the Brillouin zone. We have termed this behaviour as quasi-altermagnetism, characterized by significant momentum-dependent spin splitting with a small residual moment. Moreover, through model Hamiltonians, we investigated the microscopic origin of the quasi-altermagnetism and identified the perturbations that lead to the transition from ideal altermagnetism to quasi-altermagnetism.
@@ Acknowledgments
A.M. thanks MoE India for the PMRF fellowship. A.N. acknowledges support from Anusandhan National Research Foundation (ANRF) through grant number CRG/2023/000114. B.R.K.N. acknowledges the Department of Science and Technology (DST), India, for providing support through grant number DST/TMD/ICMAP/2K20/03(C).
Author contributions
N.D., A.N., and B.R.K.N. conceived the problem statements and designed the project. N.D. performed the DFT calculations. Symmetry analysis was carried out by A.B., N.D., A.D., A.M., and M.A.R. AHE study was performed by N.D., A.B., and M.A.R. The model Hamiltonian study was done by A.M. and B. R. K. N. All authors contributed to the analysis of the results and the preparation of the manuscript. The project was overall supervised by A. N. and B. R. K. N.
Competing interests
The authors declare no competing interests.
@@ Appendix B Symmetry operations connecting opposite spin sublattices in doped MnTe structures
For the single dopant scenario, the system consistently exhibits altermagnetic ordering. In contrast, the paired dopant scenario leads to several distinct configurations [see Fig. 3(e) of main text], with 46.67% maintaining altermagnetism, while the rest do not. Table 1 summarizes the symmetry operations connecting the opposite spin sublattices and corresponding magnetic classes for all these doped systems. For MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211), $Fm^{\prime}m^{\prime}2$ (No.42.222), and $Pm^{\prime}m^{\prime}2$ (No.25.60) the opposite spin sub-lattices are connected by symmetries other than translation or inversion, therefore being classified as altermagnets. On the other hand, for $P\bar{3}m^{\prime}1$ (No.164.89) and $C2^{\prime}/m^{\prime}2$ (No.12.62), which we refer to as quasi-altermagnets, there exist no symmetry operations that connect the up and down spin sublattices. This absence results in a lack of compensation between the spin channels, thereby lifting the degeneracy between the up and down spin bands throughout the Brillouin zone.
@@ 12.62 No symmetry Q-AM $\sigma_{xy}$, $\sigma_{yz}$
Appendix C Effect of magnetic atom doping in MnTe
The main text discusses the non-magnetic atom doping effects in altermagnetic MnTe in detail. Additionally, we have performed a preliminary investigation on the magnetic atom doping in MnTe. For this purpose, a 2 $\times$ 2 $\times$ 2 supercell of MnTe is considered and a few representative cases involving the substitution of one or two Mn atoms by Fe were analyzed.
A single substitution, as expected with different local spin moments at Fe site, breaks the rotational symmetry connecting the opposite spin sublattices, resulting in MSG $P3m^{\prime}1$ (No.156.51). The corresponding crystal structure and band structure are shown in Fig. 11 (a) and (d). The bands exhibits the characteristics of quasi-altermagnetism. In addition, new defect states are introduced, as shown in the band structure (Fig. 11 (d)).
Figure 11: Electronic band structures of Fe-doped MnTe. Panel (a) shows the MnTe supercell with a single Fe atom substituted at a magnetic site. The corresponding band structure is presented in panel (d). Panels (b) and (c) show MnTe structures doped with two Fe atoms placed at different magnetic sites, with their corresponding band structures shown in (e) and (f), respectively. The structure in (f) exhibits a quasi-altermagnetic band behavior, whereas the structure in (e) retains a perfect altermagnetic character due to the preserved mirror symmetry.
We further examined the pair substitution cases, where two Mn atoms with opposite spin moments are substituted by Fe. Similar to the pair of non-magnetic atom substitutions discussed in the main text, a pair of magnetic atom substitutions also leads to multiple configurations with different MSG. When the dopant positions are arranged in such a way that mirror symmetry (Mz) of the system is preserved, then the resulting structure belongs to MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) and will exhibit altermagnetic characteristics. The other configurations, where the proper or improper rotational symmetry is absent, lead to quasi-altermagnetism. Compared to non-magnetic atom substitution the probability of retaining altermagnetism is less when multiple magnetic atoms are substituted, since the magnetic neutrality is also affected. The other cases will result into quasi-altermagnetic conditions. Band structures displaying quasi-altermagnetic characteristics and ideal altermagnetic characteristics are shown in Fig. 11 (e) and (f). These results demonstrate that the quasi-altermagnetism discussed in the main text for non-magnetic atom doping is not restricted to such cases; it can also occur when magnetic atoms are introduced.